The general goals are to verify the flight worthiness of ACS and the functional operation of its components at temperatures beyond those expected in normal orbital conditions, and to measure its outgassing rate. The goals and procedures are described in detail in Thermal Balance/Thermal Vacuum Test Plan#3. Throughout the campaign, calibration activities will be performed by the STScI/JHU science team. These include the acquisition of dark frames for all three flight build detectors (HRC#1, WFC#4, SBC) over a range of temperatures, internal and external flat fields for the SBC to estimate count rates and characterize the delta flats, CTE measurements, and stability monitoring during transitions between hot and cold environments.
LOCATION AND DATE :
The TB/TV campaign will be conducted at GSFC over the period of May-Aug, including pre- and post-TV activities. A detailed timeline is given in the table below. ACS is in flight configuration in chamber 225 in building 7 and the science and SITS consoles and hardware are located in the nearby room 158.
INSTRUMENT CONFIGURATION :
The thermal balance fixture (TBF) holds ACS in square configuration in the vacuum chamber (the square configuration is chosen since there is no interface plate to mimic the Aft Shroud Cooling System). Photos of ACS with RAS/Cal before closing the chamber are shown in Figs 1-3. The TBF has six temperature controlled zones and with the chamber shrouds, can simulate the aft-shroud environment of HST. Diagrams of the configuration of the illumination sources for calibration purposes are shown in Figs 1 and 2. The light stimuli will be provided by a HeNe laser coupled to a single-mode fiber running through the chamber wall (see Figs 7-8) and feeding RAMP+RAS/Cal into the WFC aperture and by a PtNe backlight shining on a ~3" piece of spectralon in front of the HRC aperture. Both the PtNe lamp and the spectralon are mounted on a base plate attached to the TBF. The plate is offset so that it doesn't occult the WFC RAS/Cal beam and it (the lamp + spectralon) can be slided sideways to adjust the illumination at the HRC plane (a relatively uniform illumination is necessary for the SBC flats). The base plate can not be moved towards the instrument aperture but the lamp and spectralon can be lifted individually on the plate (with spacers under the PtNe lamp, for example). The PtNe power supply is located outside of the chamber and is nominally powered at 200 Volts and 20 mAmp for HRC coronagraphic spot tracking. The current must be adjusted appropriately for the SBC external flats to keep the global count at <3x10^5 counts/sec through different filters. We found that 14 mAmp for F165LP, 20 mAmp for F122M, and 9-10 mAmp for F115LP, are satisfactory (see Jul 18 session). Since RAMP can not be controlled remotely while in the chamber, RAS/Cal will remain aligned on the WFC for the duration of the campaign.
Before closing the chamber door, the fiber tip was aligned with the 4 um RAS/Cal pinhole and the flux of the PSF was adjusted appropriately with ND filters (see Jul 3 session). But for the first stability tracking session from TB#3 Anneal to TB#4 Hot Op, we discovered that the previous thermal transitions had offset the fiber and pinhole and that vibrations in RAS/Cal caused by the shroud blower resulted in an elongated PSF. Hence, for the Cold Soak CS#2 to Calibration transition (Jul 17) and for the orbital simulation (Jul 20), the fiber tip was illuminated and focused directly on the WFC subarray through the OPEN RAS/Cal aperture with the shroud blower turned off to minimize vibrations. The RAS/Cal tracking must be interrupted for several minutes when the temperature warms up and the blower must be turned on.
The predicted nominal detector temperatures are -77 C for WFC and -80 C for HRC. The default gain settings are 1 e-/DN for the WFC and 2 e-/DN for the HRC, with a bias offset of 3 for both detectors. The WFC will be read-out with all four amplifiers (ABCD) and the HRC with amp C. The SBC/MAMA can only be operated when the ACS internal pressure is <2x10-5 Torr. ACS will be powered on MEB side 2 for the duration of the entire campaign.
CALIBRATION ACTIVITIES :
The calibration activities are summarized in the table below as well as in sections 8.5.1-8.5.9 of the TB/TV test plan. The PtNe lamp's main purpose is to provide sufficient flux for external flats through the SBC filters/prisms in reasonable integration times and for spot tracking with HRC+Coronagraph in F606W (from the red Ne lines). Of course, the light path is not OTA-like so the low-frequency flats will be different but the high-frequency delta flats should be comparable to those acquired in the first thermal vacuum campaign (Feb 1999). Stability measurements during thermal transitions will also be performed with RAS/Cal and the HeNe point source at the WFC/F625W field centre. The measurements with the HRC+Coronagraph are the most reliable since for the WFC+RAS/Cal, the relative differences in temperature between ACS and RAS/Cal can result in spurious motions of the HeNe PSF (RAS/Cal will be moving with respect to ACS), even more so during thermal transitions, and the motions in ACS only can therefore not be isolated. The square configuration will result in severe astigmatism in the PSF but this is not important for monitoring its centroid. The use of the internal D2 lamp will be limited because of its quick degradation. We will use it to estimate count rates through the blue filters/prisms (SBC, HRC) and for some low-quality flats. The high-quality, deep flats will be obtained with the external PtNe lamp. The thorough CTE SMSs will be executed at the nominal calibration environment and detector temperatures and the CTE monitor SMSs at slightly different detector temperatures (-75 C for WFC and -83 C for HRC).
|1. Internal Calibration System Verification||HRC, WFC, SBC||T1, T2, T3, T4, D2||RT||Exercise all the internal calibration lamps, complete count rate measurements for missing filters, verify if filter in front of the SBC D2 lamp is optimal.|
|2. Basic Detector Properties||HRC, WFC, SBC||T2, T4||SMSs 01h, 01w||HRC, WFC : gain, read-noise, bias, and dark current, SBC : dark current. Measure these parameters at several temperatures besides nominal.|
|3. Detector Features||HRC, WFC||-||RT, SMSs 01h, 01w||Monitor number and location of hot pixels and columns as a function of time and temperature with dark frames.|
|4. Dark Structure||HRC, WFC, SBC||-||RT, SMSs 01h, 01w||Measure dark current structure and count rate. Also take 3600 sec SBC darks to characterize dark rate vs MAMA tube temperature.|
|5. Image Stability||HRC+Coronagraph, WFC||PtNe (HRC), RAS/Cal+HeNe (WFC)||RT (CCL loop)||Monitor the spot location on HRC or PSF centroid on WFC during thermal transitions.|
|6. Flat Field Uniformity||HRC, WFC, SBC||PtNe, Internals||RT||Verify flat field shape and structure.|
|7. Flat Field Stability||HRC, WFC, SBC||T2, T4, D2, PtNe||RT||Verify if flats change shape with time and for the CCDs, with the detector temperature. Compare with deep internal flats (F435W, F625W, and F814W) of Mar and Apr 2001.|
|8. SBC Flats||SBC||PtNe, D2||RT||Acquire a set of deep external PtNe flats at ~10 ke-/pix in F125LP and F165LP. Also acquire a deep internal D2 flat through F125LP and shorter exposures through the remaining filters/prisms to estimate the internal count rates.|
|9. Charge Transfer Efficiency (CTE)||HRC, WFC||T2, T4||SMSs 02h, 02w, 05h, 05w||Perform thorough CTE measurements (05h, 05w) at the nominal detector temperatures and monitor (02h, 02w) at other temperatures.|
The campaign consists of a Thermal Balance (TB) session followed by a Thermal Vacuum (TV) session. In the TB program, ACS will be subjected to four different orbital environments to be encountered in the HST aft shroud (TB#1, #2, #3, and #4). The TV session includes two worse case cold (cold soaks #1, #2) and hot (hot soaks #1, #2) environments at -32 C and +4 C, respectively, approximately 10 C more severe than the expected orbital temperatures. The bulk of the science calibration activities are scheduled in the last ~5 days of TV at a nominal aft shroud temperature of -8 C environment but some (such as image stability) will be executed whenever the schedule allows. In the following table, we list the test procedure of the campaign as given in the TB/TV plan and its timeline, and the possible calibration activities associated with each. Of course, the scheduling of the activities is subject to change as the campaign progresses.
|9.1 Installation of Test Equipment in the TV Chamber for Pre-Test Certification|
|188.8.131.52 - 184.108.40.206||May 16-23||-||-|
|9.2 Pre-Test GSE/Chamber Outgassing Certification|
|220.127.116.11 - 18.104.22.168||May 24 - Jun 18||-||-|
|9.3 Installation of ACS and the Test Equipment in TV Chamber 225|
|22.214.171.124 - 126.96.36.199||Jun 16-29||-||-|
|9.4 Thermal Balance|
|188.8.131.52 - 184.108.40.206
|Jul 2-3||-||Pre-TV3 check-out of RAS/Cal and PtNe illumination|
|220.127.116.11 - 18.104.22.168
|Jul 5 (09:00)-6 (11:00)||-||TBF heater circuits set at +22 C|
|22.214.171.124 - 126.96.36.199
Cold Safe : Thermal Balance #1
|Jul 6 (12:10) - 7 (13:00)||-||Jul 6 (11:15) : 188.8.131.52 : Perform AT on MEB sides 1
184.108.40.206 : Turn ACS on and take test exposures with RAS/Cal and PtNe. We find that thermal effects moved the RAS/Cal spot by ~30 pix in X and Y relative to pre-TV3 check-out and it went out of focus.
|220.127.116.11 - 18.104.22.168
Cold Operation : Thermal Balance #2
|Jul 7 (16:40)-8 (11:30)||-||Jul 8 (12:30) : 22.214.171.124 : Turn TECs on and perform HRC and WFC temperature walk-downs in this cold operate environment.|
|126.96.36.199 - 188.8.131.52
Cold Anneal : Thermal Balance #3
|Jul 8 (17:10)-9 (14:30)||-||Jul 9 (15:26) : 184.108.40.206 : ACS in WFC/HRC Operate Mode|
|220.127.116.11 - 18.104.22.168
Hot Operation : Thermal Balance #4
|Jul 9 (20:30)-11 (03:00)||5||22.214.171.124-126.96.36.199 : Image stability during transition
from TB#3 Anneal to TB#4 Hot Op : TECs on (HRC : -81.2 C, WFC : -76.6
C), coronagraph arm deployed, RAS/Cal, HeNe, PtNe on.
(1) Jul 9 (16:00-20:00) : Attempt to get a RAS/Cal HeNe point source on WFC at end of TB#3 Anneal. Two problems : (1) pinhole and fiber tip are not aligned (2) vibration in RAS/Cal from the shroud blower being turned on results in an elongated PSF (see ID 30201).
(2) Jul 9 (20:00)-11 (03:00) : HRC+PtNe coronagraph spot tracking only (see ID 30202) : Jul 9 (20:00-20:30) : 1/2 hour baseline, Jul 9 (20:30)-11 (03:00) : Transition from TB#3 to TB#4. The results are available.
(3) Jul 11 (10:00) : 188.8.131.52-184.108.40.206 : Temperature walk-down, D2 on/off, WFC tungsten on/off
|220.127.116.11 - 18.104.22.168
MAMA "ON" sensitivity at hot operation thermal balance #4
Jul 12 (09:45) : Stabilization at WFC/MAMA hot condition is declared.
22.214.171.124 : Acquire SBC darks (~1000 sec) as the MAMA tube warms up from its nominal temperature in RT.
Note : Jul 11-12 : An internal ACS pressure of <2x10-5 Torr could not be reached so this item was skipped.
|9.5 Thermal Vacuum|
|126.96.36.199 - 188.8.131.52
Transition to hot soak #1
|Jul 12 (09:45-10:00)||-||ACS in WFC/HRC Operate Mode|
|184.108.40.206 - 220.127.116.11
Hot soak #1
|Jul 12 (10:00-13:00)||-||-|
|18.104.22.168 - 22.214.171.124
Transition to cold soak #1
|Jul 12 (13:00)-Jul 12 (23:54)||-||Jul 12 (23:54) : 126.96.36.199 : ACS in OFF mode for 1 hr|
|188.8.131.52 - 184.108.40.206
Cold soak #1 (with cold start on side 1)
|Jul 13 (00:54-17:15)||2||(a) Jul 13 (00:54) : 220.127.116.11 : ACS is turned on
in WFC/HRC Operate Mode, TECs at nominal setpoints from 18.104.22.168
(b) Jul 13 (08:30-13:00) : 22.214.171.124 Execute basic CCD monitor SMSs
(c) Jul 13 (13:00-16:30) : 126.96.36.199 Mech. relay test
(d) Jul 13 (16:30-17:15) : TECs on until HRC is at -80 C and WFC at -77 C.
|188.8.131.52 - 184.108.40.206
Transition to Hot Soak #2, Functional Test
|Jul 13 (17:15)-Jul 14 (00:45)||5||Image stability during transition from Cold Soak #1 to Hot Soak #2 : TECs on, coronagraph arm deployed, PtNe on. Note : No RAS/Cal+HeNe point source tracking. Jul 13 (17:15-18:00) : Baseline, Jul 13 (18:00)-Jul 14 (00:45) : Transition. The results are available.|
|220.127.116.11 - 18.104.22.168
Hot Soak #2, Functional Test
|Jul 14 (00:45)-Jul 15 (13:05)||-||Imaging and servicing functionals on MEB sides 1 and 2.
22.214.171.124 : Jul 14 (00:58) : ACS safed during voltage testing
|126.96.36.199 - 188.8.131.52
Transition to Cold Soak #2, Functional Test
|Jul 15 (13:05-22:20)||-||Jul 15 (23:20) : 184.108.40.206 : ACS in OFF mode for 1 hr|
|220.127.116.11 - 18.104.22.168
Cold Soak #2, Functional Test (with cold start on side 2)
|Jul 15 (22:20)-Jul 17 (14:50)||-|| (a) ACS is turned on and in WFC/HRC Operate
Mode. Internal pressure at ~1x10-5 Torr.
(b) Imaging and servicing functionals on MEB sides 1 and 2.
|22.214.171.124 - 126.96.36.199
Transition from CS#2 to Calibration
|Jul 17 (14:50-21:45)||5||(a) Jul 17 (10:00-14:50) : Managed to illuminate and
focus the fiber tip directly on a WFC subarray through the OPEN
RAS/Cal aperture by turning off the shroud blower to minimize
vibrations. The resultant point source (example : ID 31873) suffers significant
astigmatism. This configuration will be used for point source
tracking in the following transition.
(b) Jul 17 (14:50-21:45) : Transition : Stability monitoring with both HRC+Coronagraph (PtNe at 20 mAmp) and RAS/Cal+HeNe. Note : Shroud blower was turned on twice so RAS/Cal tracking was turned off during those periods. The results of the HRC+Coronagraph spot tracking and the WFC+RAS/Cal point source tracking are available.
|188.8.131.52 - 184.108.40.206
|Jul 17 (22:00)-Jul 21 (20:00)||1, 2, 3, 4, 5, 6, 7, 8, 9||
(a) Jul 17 (22:00)-Jul 18 (09:00) : 220.127.116.11 : HRC and WFC basic
monitor SMSs (bias, darks, internal flats) on MEB side 2 at all
supported gain settings followed by WFC darks at -77.3 C, -76.3 C, and
-75 C. Internal ACS pressure at 2x10-5 Torr.
(b) Jul 18 (09:00-11:00) : 18.104.22.168 : Use SBC "safe-open" procedure, then acquire internal D2 flats on SBC through all filter wheel 3 filters/prisms to estimate count rates with the ND1.3 filter.
(c) Jul 18 (11:00-16:45) : 22.214.171.124 : Use SBC "safe-open" procedure, then acquire a series of 1000 sec SBC delta flats through F115LP (instead of F125LP, as quoted in calibration item 8) with the internal D2 lamp at levels of 10 ke- per 4 pix resolution element.
(d) Jul 18 (17:20)-Jul 19 (16:00) : 126.96.36.199 : Use SBC "safe-open" procedure, then acquire high S/N-ratio SBC flats through F115LP (instead of F125LP, as quoted in calibration item 8 above) with the external PtNe backlight to levels of 10 ke-/pix. The dataset consists of 10x3600 sec exposures followed by 4x3600 sec SBC darks. A test exposure through F165LP was also acquired followed by a sequence of 3x3600 sec external PtNe flats (the "paw" print pattern is still visible). Similarly, 3x3600 sec external flats were acquired through F122M.
(e) Jul 19 (16:00)- Jul 20 (16:30) : 188.8.131.52 : Execute HRC and WFC thorough CTE SMSs (note : only FPR portion of JTVW05B (amps AD) was executed, to be resumed later) and acquire associated biases and darks for reduction. Obtained a suite of bias frames for delivery to STScI.
(f) Jul 20 (16:30-21:00) : Unplanned calibration item : Stability monitoring with HRC+Coronagraph (PtNe lamp) and WFC+RAS/Cal (HeNe) fiber point source. The -V2, -V3, and aft panels are warmed up and cooled down by 8 C every ~45 min from their nominal starting temperatures to simulate the orbital day/night thermal transitions. The HRC+Coronagraph and WFC+RAS/Cal stability curves are available.
(g) Jul 20 (21:20)-Jul 21 (05:00) : Complete last FPR image and EPER portion of JTVW05B followed by the CTE monitor SMSs JTVW02A and JTVH02A at -75 C for the WFC and at -83 C for the HRC.
(h) Jul 21 (05:00-11:45) : Investigation of the HRC column anomaly (two datasets), suite of bias frames for pipeline delivery, coherent noise on HRC at low bus voltage as seen in CS#2 during the SMFT.
(i) Jul 21 (11:45-18:55) : 184.108.40.206 : Acquire a set of 1000 sec darks and 1000 sec F115LP external flats (PtNe illumination) as a function of increasing MAMA tube temperature. Perform this item only if 220.127.116.11 fails, otherwise skip.
(j) Jul 21 (18:55-20:00) : Re-visit coherent noise on HRC with HRC/CEB cold.
|18.104.22.168 - 22.214.171.124
Transition to the Outgassing Certification
|Jul 21 (20:00)-Jul 22 (02:00)||-||-|
|126.96.36.199 - 188.8.131.52
|9.6 Transition to Ambient Temperature|
|184.108.40.206 - 220.127.116.11||Jul 22 (14:00)-Jul 23 (02:00)||-||-|
|9.7 Chamber Backfill, Ambient Abbreviated Funtional Test|
|18.104.22.168 - 22.214.171.124||-||-||-|
|9.8 End of Test|